M7_IND6. Cavendish Convenience Stores purchases 1000 U.S. Road Maps a year for its inventory from
its supplier, who offers pricing at quantity discounts. The cost for Cavendish to place an order is $42, and
the annual carrying cost is 15%/year. What quantity should Cavendish order? The quantities and pricing
from this supplier are shown in the following table:
ORDER QUANTITY UNIT PRICE
0-149 $12.00
150-349 $11.70
350-999 $11.50
1000 or more $11.20
a) What is the EOQ when the unit price is $12?
b) What is the total cost (holding + purchase+ ordering) when purchasing from the supplier where
the unit price is $11.70?
c) What is the total cost (holding + purchase+ ordering) when purchasing from the supplier where
the unit price is $11.50?
d) What is the total cost (holding + purchase+ ordering) when purchasing from the supplier where
the unit price is $11.20?
e) Based on the total costs, what quantity should be ordered from the supplier?
4

Show work
Answer #1

A. D= 1000 = yearly demand

S = ordering cost = 42 dollars

i = 0.15

C=unit price = 12

/2DS

= 216.02>150 so EOQ = 150

Total cost = 150*0.15*12/2 + 1000/150*42 + 1000*12 = 12415 dollars

B. C= 11.70

/2DS

=218.78

Total cost = 218.78*0.15*11.7/2 + 1000/218.78*42 + 1000*11.7 = 12083.95 dollars

C. C= 11.5

/2DS

=220.67< 350

So EOQ =350

Total cost = 350*0.15*11.5/2 + 1000/350*42 + 1000*11.5 = 11921.875 dollars

D. C =11.2

/2DS

=223.61<1000

So EOQ = 1000

Total cost = 1000*0.15*11.2/2 + 1000/1000*42 + 1000*11.2 = 12082 dollars

D. min cost is 11921.875 dollars at EOQ = 350 units